### v1

# Cooper quartets designing in multi-terminal superconducting devices

Cooper quartets designing in multi-terminal superconducting devices
Luca Chirolli NEST, Istituto Nanoscienze - CNR, and Scuola Normale Superiore, I-56127 Pisa, Italy    Alessandro Braggio NEST, Istituto Nanoscienze - CNR, and Scuola Normale Superiore, I-56127 Pisa, Italy    Francesco Giazotto NEST, Istituto Nanoscienze - CNR, and Scuola Normale Superiore, I-56127 Pisa, Italy
Quantum design of Cooper quartets in a double quantum dot system coupled to ordinary superconducting leads is presented as a novel platform for the study of an elusive many-body state of matter, that is at the basis of the phenomenon of charge-$4e$ superconductivity. A fundamentally novel, maximally correlated ground state, in the form of a superposition of vacuum $|0\rangle$ and four-electron state $|4e\rangle$, emerges as a narrow resonance and it is promoted by an attractive interdot interaction. A novel phenomenology in the dissipationless transport regime is elucidated, that yields typical flux quantization in units of $h/4e$ and manifests in non-local multi-terminal coherence and in two-Cooper pair transport properties mediated by the quartet ground state. The results open the way to the exploration of correlation effects and non-local coherence in hybrid superconducting devices, parity-protected quantum computing schemes and more generally, the work poses the basis for the design and simulation of novel correlated states of matter starting from ordinary ingredients available in a quantum solid state lab.
The Cooper instability predicts a state of matter in which electrons pair up and form a condensate showing unique properties such as zero resistivity and perfect diamagnetism. Making a conceptual jump we can ask ourselves if more complicate electron aggregates such as Cooper quartets could form and condense in a novel state that gives rise to charge-$4e$ superconductivity. In spin 1/2 systems, Pauli’s exclusion principle forbids local multiparticle aggregates beyond pairs and a Bardeen-Cooper-Schrieffer superconductor constitutes a weakly correlated condensate of pairs in which quartet correlations, that are absent in the mean-field ground state, may emerge only as a residual of interaction among pairs. Purely four-fermion instabilities have been predicted in higher spin systems Wu (2005) or in systems showing an additional quantum number, such as $\alpha$-particle in nuclear physics Röpke et al. (1998). It has been suggested that a charge-$4e$ superconducting state may emerge in the fluctuating state, where Cooper pairs are loosely bound, and in systems displaying a two-component condensate Berg et al. (2009); Lee (2014); Fradkin et al. (2015), such as copper-oxides Li et al. (2007); Ding et al. (2008), kagome systems Lin and Nandkishore (2021); Chen et al. (2021); Zhou and Wang (2022); Li et al. (2023); Guguchia et al. (2023); Yu (2023), twisted bilayer graphene Jian et al. (2021); Fernandes and Fu (2021); Liu et al. (2023), iron-based superconductors Fernandes et al. (2014), or doped topological insulators Matano et al. (2016); Shen et al. (2017). Effective charge-$4e$ superconducting transport properties are obtained in networks of ordinary Josephson junctions, where transport of Cooper pairs is suppressed by destructive interference effects and only two-Cooper pair transport can survive, yielding a phenomenology similar to the one associated to the proper charge-$4e$ state Douçot and Vidal (2002); Douçot et al. (2003); Protopopov and Feigel’man (2004); Douçot et al. (2005); Rizzi et al. (2006).
In this work, we change the perspective and resort to the art of quantum design to tailor the conditions for the onset of a Cooper quartet ground state. First of all, in order to fully study the emergence of many-body correlations we consider a minimal model of a double quantum dot, that allows us to accommodate a four-electron superposition state through an additional orbital binary quantum number beyond spin. In addition, we exploit Cooper pairing as already provided by nature through proximity coupling the double-dot system to a ordinary superconductor. Finally, in order to stabilize the quartet phase we suggest a mechanism based on attractive interdot interaction, that despite its unusual character has been experimentally realized through a second double quantum dot Hamo et al. (2016), or through a transmission line resonator Delbecq et al. (2013). The system manifests a quartet ground state
$|\phi^{\pm}_{\rm Q}\rangle=\frac{1}{\sqrt{2}}(|0\rangle\pm|4e\rangle),$ (1)
in which the vacuum state $|0\rangle$ and the four-electron state $|4e\rangle$ appear in superposition of equal weights and that describes the formation of a coherent, albeit minimal, condensate, whose properties are dictated by the charge $4e$. We show that many-body interactions among the dots yield a finite purely quartet correlator,
$Q$ $=$ $\langle d_{1\downarrow}d_{1\uparrow}d_{2\downarrow}d_{2\uparrow}\rangle-\langle d_{1\downarrow}d_{1\uparrow}\rangle\langle d_{2\downarrow}d_{2\uparrow}\rangle$ (2)
-- $\langle d_{1\downarrow}d_{2\uparrow}\rangle\langle d_{1\uparrow}d_{2\downarrow}\rangle,$
with $d_{i\sigma}$ fermionic dot annihilation operators, showing that a quartet scattering amplitude has a purely interacting origin. The correlator $Q$ resonantly peaks to its extreme values ±$\pm 1/2$ when the states $|\phi_{Q}^{\pm}\rangle$ are realized. A manifestation of the charge $4e$ nature of the ground state appears when ordinary superconducting leads are attached to the double-dot system and a charge-$4e$ Josephson effect appears. Upon closing the double-dot system on a superconducting loop threaded by a magnetic flux, the ground state current yields a flux quantization in units of $h/4e$. The quartet ground state delocalized in the two quantum dots shows a strong non-local coherent character and we identify a novel non-local phase response in a three-terminal setup that allows us to harvest the correlated nature of the ground state.
Figure 1: a) Schematics of the double-dot system with onsite interaction $U$, interdot density-density interaction $W$, tunnel-coupled to a superconductor $S$ through rates $\gamma_{S}$ and $\gamma_{S}^{\prime}$. b) Scheme of energy levels and their coupling when the resonance condition is met, $\epsilon=-U/2-W+\delta\epsilon$, up to a small detuning $\delta\epsilon$. c) Phase diagram of the quartet correlator $Q$ at resonance $\epsilon=-U/2-W$ for $\gamma_{S}^{\prime}=\gamma_{S}$: for $-W>U>0$ the ground state has a value $Q=-1/2$ compatible with a quartet state $|\phi^{-}_{Q}\rangle$, whereas for $U<W<0$ the ground state has a value $Q=1/2$ associated to the quartet state $|\phi^{+}_{Q}\rangle$. d) Quartet resonance as a function of $\delta\epsilon=\epsilon+U/2+W$ along the dashed gray line in c), where $U-W=3\gamma_{S}$, showing the dependence of the correlator $Q$ on the detuning of the vacuum and four-electron states.
These findings open the way to interaction effects in superconductors, where non-local phase coherence can be establish in multi-terminal setups, and it can find application in parity-protected quantum computing schemes, such as those based on two-Cooper pair transport Blatter et al. (2001); Ioffe et al. (2002); Ioffe and Feigel’man (2002); Brooks et al. (2013); Douçot and Ioffe (2012). More generally, the results presented open the way to the design and simulation of novel phases of matter based on exotic electron complexes. Through manipulation of electron interactions in few body systems with tools and ingredients available in a solid-state quantum lab, it is possible to study the properties of otherwise inaccessible states of matter.
The model.—We consider the double quantum dot system, with the two dots labeled with $i=1,2$, tunnel coupled to a common superconducting lead $S$ at the left of the system, as schematized in Fig. 1a). We describe their general Hamiltonian $H_{0}$ through single-electron gate-tunable levels with energy $\epsilon_{i}$, an intradot generally repulsive Hubbard term with strength $U$, and an interdot density-density interaction with strength $W$, that we specifically allow to be negative, so to describe an attraction between the dots Hamo et al. (2016); Delbecq et al. (2013). The dot Hamiltonian then reads
$H_{0}$ $=$ $\sum_{i,\sigma}\epsilon_{i}n_{i\sigma}+U\sum_{i}n_{i\uparrow}n_{i\downarrow}+Wn_{1}n_{2},$ (3)
with $n_{i}=n_{i\uparrow}+n_{i\downarrow}$, $n_{i\sigma}=d^{\dagger}_{i\sigma}d_{i\sigma}$, and $d_{i\sigma}$ dot fermionic annihilation operators. We then describe tunnel coupling to the superconducting lead $S$ as a proximity effect, for which the quantum dots develop a finite pairing amplitude to form Cooper pairs, either locally on each dot or in a delocalized way, between the two dots, so that the pairing Hamiltonian reads
$H_{p}=\gamma_{S}\sum_{i}d^{\dagger}_{i\uparrow}d^{\dagger}_{i\downarrow}+\gamma^{\prime}_{S}(d^{\dagger}_{1\uparrow}d^{\dagger}_{2\downarrow}-{d^{\dagger}_{1\downarrow}d^{\dagger}_{2,\uparrow}})+{\rm H.c.}.$ (4)
This model has been thoroughly studied in the limit in which the scattering off the superconducting lead becomes elastic and integration away of the leads in the $\Delta\to\infty$ limit produces effective pairing amplitude, $\gamma_{S}=2\pi\nu_{F}t^{2}_{DS}$ and $\gamma^{\prime}_{S}=\frac{1}{2}\gamma_{S}e^{-\delta r/\xi}\sin(k_{F}\delta r)/(k_{F}\delta r)$, that coincides with the rates of local and crossed Andreev reflection, with $t_{DS}$ the dot-superconductor tunneling amplitude, assumed equal for the two dots, $\delta r$ the interdot distance, $k_{F}$ the Fermi momentum, and $\xi$ the coherence length of the superconductor. The full Hamiltonian is $H=H_{0}+H_{p}$ and we focus the analysis on the even parity sector, where the four electron operator $d_{1\downarrow}d_{1\uparrow}d_{2\downarrow}d_{2\uparrow}$ can have non-zero matrix element.
Quartet correlations emerge specifically when the resonance $04\epsilon+2U+4W=$ 0 occurs, for which the vacuum $|0\rangle$ and the four-fold occupied state $|4e\rangle$ become degenerate. To further select them as a host for a quartet ground state, we require all other states to have higher energy and, assuming a repulsive onsite interaction $U>0$, a relevant condition is met when the quantum dots experience an attractive density-density interaction that is stronger than the onsite repulsion, $-W>U>0$. At resonance, $\epsilon=-U/2-W>0$, and for $-2W>0$, $-U-W>0$, and $\gamma_{S},\gamma_{S}^{\prime}\ll|U+W|,U,|W|$, the vacuum and the four-electron states couple at first order with the higher energy states $|\phi_{2e}^{+}\rangle\equiv(|{\uparrow\downarrow}\rangle_{1}+|{\uparrow\downarrow}\rangle_{2})/\sqrt{2}$ and $|\phi_{2e}^{-}\rangle\equiv(|{\uparrow}\rangle_{1}|{\downarrow}\rangle_{2}-|{\downarrow}\rangle_{1}|{\uparrow}\rangle_{2})/\sqrt{2}$, as schematized in Fig. 1b) and detailed in  [^1] Supplemental Material containing details on the Hamiltonian and the correlator., and results in the low energy states
$|\bar{0}\rangle=|0\rangle+\frac{\sqrt{2}\gamma_{S}}{2W}|\phi^{+}_{2e}\rangle+\frac{\sqrt{2}\gamma^{\prime}_{S}}{U+W}|\phi^{-}_{2e}\rangle,$ (5)
$|\bar{4e}\rangle=|4e\rangle+\frac{\sqrt{2}\gamma_{S}}{2W}|\phi^{+}_{2e}\rangle-\frac{\sqrt{2}\gamma^{\prime}_{S}}{U+W}|\phi^{-}_{2e}\rangle.$ (6)
We then project the Hamiltonian on the low energy states, where it assumes a simple form
$h=-2\delta\epsilon(|\bar{0}\rangle\langle{0}|-|\bar{4e}\rangle\langle\bar{4e}|)+\Gamma(|\bar{0}\rangle\langle\bar{4e}|+{\rm H.c.}),$ (7)
with the coupling matrix element $\Gamma=-\frac{2(\gamma_{S}^{\prime})^{2}}{U+W}+\frac{\gamma_{S}^{2}}{W}$ and $\delta\epsilon$ a weak detuning from resonance, $\epsilon=-U/2-W+\delta\epsilon$.
Figure 2: a) Schematics of the three-terminal structure composed by two quantum dots, 1 and 2, coupled to a common superconducting lead $S_{0}$ at the left and to superconducting leads, $S_{1}$ and $S_{2}$, at the right. The latter are biased at phase $\varphi_{1}$ and $\varphi_{2}$ with respect to $S_{0}$, respectively. b) Contour plot of the ground state energy of the system in a) with respect to the phases $\varphi_{1}$ and $\varphi_{2}$, showing two minima at approximately $\varphi_{1}=\varphi_{2}\simeq\pm\pi/2$, for $\gamma_{S}^{\prime}=\gamma_{S,l}=\gamma_{S,r}=\gamma_{S}$, $\epsilon=-U/2-W$, $U/\gamma_{S}=10$ and $W/\gamma_{S}=-12$. c) Total current $I_{1}+I_{2}$ as a function of $\varphi_{1}=\varphi_{2}$ along the white line in b), shown for the case $\gamma_{S,l}=\gamma_{S,r}$ as in b) and for $\gamma_{S,l}=0$ for which the current is exactly $\pi$-periodic.
The phase diagram of the system is shown in Fig. 1c), where we plot the ground state value of correlator $Q$ at resonance $\epsilon=-U/2-W$, for positive and negative values of $U$ and $W$. The correlator can take generic complex values such that $|Q|\leq 1/2$ and for real $\gamma_{S},\gamma_{S}^{\prime}$ it is real. In the region $U>W$ the ground state belongs to the subspace spanned by $\{|\phi^{-}_{Q}\rangle,|\phi^{-}_{2e}\rangle\}$ and the correlator evolves from $Q=-1/2$, when $U+W<0$ and the ground state has a strong $|\phi^{-}_{Q}\rangle$ quartet character, to $Q\sim(\gamma_{S}^{\prime})^{2}/(U+W)^{2}$ when $U+W>0$, where the ground state has mostly a two-electron state character $|\phi^{-}_{2e}\rangle$. In the region $U<W$ of the phase diagram Fig. 1c) the ground state belongs to the sector spanned by $\{|\phi^{+}_{Q}\rangle,|\phi^{+}_{2e}\rangle\}$, and for the fully attractive case $W,U<0$ a quartet ground state $|\phi^{+}_{Q}\rangle$ is achieved for which $Q=1/2$. Moving away from resonance, for $\epsilon=-U/2-W+\delta\epsilon$, the value of the correlator gives a Breit-Wigner resonance $Q\simeq-\Gamma/(2\sqrt{\Gamma^{2}+4\delta\epsilon^{2}})$ with linewidth $\Gamma$ for $U+W<0$, as shown in Fig. 1d), and away from resonance the quartet correlations rapidly decaying.
Dissipationless transport.— It becomes natural at this point to search for a direct consequence of having a quartet ground state and a smoking gun observable that can unambiguously witness its presence. We first look at the equilibrium properties and dissipationless transport in a three-terminal structure constituted by a common lead at the left, relabeled lead 0, and two additional leads 1 and 2 each coupled to the corresponding quantum dot, as schemitized in Fig. 2a), and study the supercurrents in the system. Proximity induced local pairing in quantum dot $i=1,2$ can originate either from lead 0 or from the corresponding lead, 1 or 2, at the right, biased at phase $\varphi_{i}$ with respect to lead 0, as schematized in Fig. 2 a). We assume different local rates at the left $\gamma_{S,l}$ and at the right $\gamma_{S,r}$ leads and equal for the two dots. This way, local pairing in dot $i$ becomes controlled by $\gamma_{S,i}=\gamma_{S,l}+\gamma_{S,r}e^{i\varphi_{i}}$. In turn, non-local pairing $\gamma_{S}^{\prime}$ can only result from Cooper pair splitting from lead 0 and it is not sensitive to any phase differences. The Hamiltonian projected on the low energy manifold reads
$h$ $=$ $\frac{\gamma_{S}^{2}}{W}(2+\cos(\varphi_{1})+\cos(\varphi_{2}))$ (8)
-- $2\delta\epsilon(|\bar{0}\rangle\langle\bar{0}|-|\bar{4e}\rangle\langle\bar{4e}|)+\Gamma|\bar{4e}\rangle\langle\bar{0}|+{\rm H.c.},$
that generalizes the coupling between vacuum and four-electron states to a complex one that, for $\gamma_{S,l}=\gamma_{S,r}=\gamma_{S}$, reads
$\Gamma(\varphi_{1},\varphi_{2})=-\frac{2(\gamma_{S}^{\prime})^{2}}{U+W}+\frac{\gamma^{2}_{S}}{W}(1+e^{i\varphi_{1}})(1+e^{i\varphi_{2}}).$ (9)
Its argument $\Theta={\rm arg}\left[\Gamma\right]$ represents the phase of the superposition between vacuum and four-electron state in the ground state, $|\Psi_{\rm GS}\rangle=\frac{1}{\sqrt{2}}(|\bar{0}\rangle-e^{i\Theta}|\bar{4e}\rangle)$, and can be tuned through the phases $\varphi_{1}$ and $\varphi_{2}$.
Figure 3: a) Current $I_{1}(\varphi_{1},\varphi_{2}=0)$ and b) $I_{2}(\varphi_{1},\varphi_{2}=0)$ as a function of the phase $\varphi_{1}$ and the detuning $\delta\epsilon$, obtained by keeping $\varphi_{2}=0$. c) Cuts of $I_{1}(\varphi_{1},\varphi_{2}=0)$ in a) at three different values of $\delta\epsilon$ shown in the legend of d). d) Cuts of $I_{2}(\varphi_{1},\varphi_{2}=0)$ at the values of $\delta\epsilon$ specified in the legend.
The most intuitive property a quartet ground state is expected to show is to support a two-Cooper pair current between externally attached superconducting leads. At an ordinary Josephson junction the tunneling of a Cooper pair amounts to a $2\pi$ slip in the phase difference at the junction, so that the coherent tunneling of two Cooper pairs yields a $\pi$-periodic Josephson current. The ground state energy is shown in Fig. 2b) and it clearly shows two minima at $\varphi_{1}=\varphi_{2}\simeq\pm\pi/2$. The equality does not hold in a strict way as a consequence of weak but finite single Cooper pair contribution in the lowest energy states $|\bar{0}\rangle$ and $|\bar{4e}\rangle$. Associated to the quartet ground state, we expect the system to support a quartet current, so that we join contact 1 and 2 in a unique drain by setting $\varphi_{1}=\varphi_{2}=\varphi$ and study the total current $I=I_{1}+I_{2}$. The ground-state current through dot $i$ features two contributions: an ordinary Josephson term describing current between leads $i$ and 0, and a second term that depends on both $\varphi_{1}$ and $\varphi_{2}$,
$I_{i}=-\frac{2e}{\hbar}\frac{\gamma_{S}^{2}}{W}\sin(\varphi_{i})-\frac{2e}{\hbar}\frac{\partial}{\partial\varphi_{i}}\left|\Gamma(\varphi_{1},\varphi_{2})\right|.$ (10)
It is clear that a $\pi$-periodic current can only arise approximately, in that the contribution coming from $|\Gamma|$ must cancel the ordinary $2\pi$-periodic pair contribution of the first term. Assuming $W=-U-\delta W$, $\delta W\ll U$, $\gamma_{S}^{\prime}=\gamma_{S}$, and expanding the current at lowest orders in $\delta W/U$ we obtain
$I=-\frac{4e\gamma^{2}_{S}}{\hbar U}\left[\sin(2\varphi)-{\cal O}\left(\delta W/U\right)\right].$ (11)
The cancellation of the first harmonic is possible through a destructive interference effect originating from the opposite sign in the expression of $\Gamma$ between the phase-independent term describing non-local Cooper pair splitting and the phase-dependent term resulting from local Cooper pair tunneling. The result can be understood as follows: if we could ideally select only Cooper pair splitting at the left lead, by setting the local contribution to zero $\gamma_{S,l}=0$, and only direct Cooper pair tunneling at the right—as the structure with two terminals naturally selects—the only possible process would be a two-Cooper pair current, as shown in Fig. 2c) for comparison and explained in Note (1). For finite $\gamma_{S,l}=\gamma_{S,r}=\gamma_{S}$, we can obtain it approximately by promoting the relative importance of the non-local process with respect to the local one in the expression of $\Gamma$ by requiring $|2(\gamma_{S}^{\prime})^{2}/(U+W)|\gg|\gamma_{S}^{2}/W|$ so to effectively suppress the single Cooper pair current. The resulting current is shown in Fig. 2c) and shows a dominant $-\sin(2\varphi)$ component. Importantly, we obtain such a condition without recurrence to any flux interferometric methods Douçot and Ioffe (2012).
The interacting nature of the quartet correlator suggests us that another possible clear signature of the quartet ground state can appear in the non-local response of the system. Indeed, a given phase difference established between two given terminals affects the current through the third terminal in a non-trivial way. If we close lead 2 on lead 0 and enclose a flux $\Phi_{x}$ we effectively pin the phase $\varphi_{2}=2e\Phi_{x}/\hbar$, so that we can selectively choose an horizontal cut in the energy landscape Fig. 2b) and study the associated currents in terminal 1 and 2. The resulting $I_{1}$ is shown in Fig. 3a) where application of a phase bias $\varphi_{1}$ between terminal 1 and 0 at constant $\varphi_{2}=0$ yields a sinusoidal current $I_{1}(\varphi_{1})$ away from the quartet resonance condition that sharply evolves in an opposite current at resonance $\delta\epsilon=0$, in an effective $\pi$-junction behavior. The latter is visualized in Fig. 3c) for three values of $\delta\epsilon$. Even more counter-intuitive is the current $I_{2}(\varphi_{1})$ that flows in the ring at zero external flux. $I_{2}$ activates only at resonance and shows a finite signal concord to $I_{1}$, as shown in Fig. 3b) and d).
Conclusions.— We have demonstrated how appropriate tuning of electron-electron interaction in a double quantum dot system opens the possibility to design quartet superconductivity. The system presented shows a great host of counter-intuitive properties and shall open unexplored paths in the simulation of correlated states of matter that hardly occur in nature, from charge-$4e$ superconductivity to more exotic phases based on novel electron complexes. Once a solid quartet ground state is stabilized a first phenomenological expectation is flux quantization in units of $h/4e$. To this end, contacts $S_{1}$ and $S_{2}$ need to be joined in a right superconducting lead, with care to not generate non-local Andreev reflection at the right lead, as explained in Note (1). Upon closing the circuit on a superconducting loop enclosing a flux $\Phi_{x}$, a charge $4e$ current appear as the $\pi$-periodic current in Fig. 2c), with $\varphi=2e\Phi_{x}/\hbar$. A consequent very appealing application is to design a parity-protected qubit, where two-Cooper pair transport yields quantum states with a definite parity of the number of Cooper pairs Blatter et al. (2001); Ioffe et al. (2002); Ioffe and Feigel’man (2002); Brooks et al. (2013); Douçot and Ioffe (2012). The possibility to experimentally achieve the isolation of a quartet ground state relies on the ability to engineer an attractive density-density interaction, that can be mediated by a second double quantum dot that is capacitively coupled to the system Hamo et al. (2016), or through a transmission line resonator Delbecq et al. (2013). Our results stimulate further research to consolidate the platform and holds promises for the design and simulation of highly correlated phases of matter based on complex electron aggregates.
Acknowledgment.—We acknowledge the EU’s Horizon 2020 Research and Innovation Framework Programme under Grant No. 964398 (SUPERGATE), No. 101057977 (SPECTRUM), and the PNRR MUR project PE0000023-NQSTI for partial financial support. A.B. acknowledges the Royal Society through the International Exchanges between the UK and Italy (Grants No. IEC R2 192166.)